OpenStudy (babynini):

Solve for theta y=cos(theta/2)

OpenStudy (babynini):

\[\theta=2\cos^{-1}(y)\]

OpenStudy (anonymous):

They are the same thing, so it's true for all angles theta

OpenStudy (babynini):

The actual question itself has more components. \[x=\sin \frac{ 1 }{ 2 }\theta, y=\cos \frac{ 1 }{ 2 }\theta\] I want to write it either in terms of y or x \[\theta=2\cos^{-1}(y) \]\[x=\sin \frac{ 1 }{ 2 }*2\cos^{-1}(y)\]\[x=\sin(\cos^{-1}(y))\]

OpenStudy (babynini):

@larryboxaplenty what is true? o_o

OpenStudy (babynini):

Once we know theta can't we just plug it into x and simplify?

OpenStudy (babynini):

The question: "Eliminate the parameter to find a Cartesian equation of the curve" \[-\pi \le \theta \le \pi\] then look up a few comments for the rest of the question xD

OpenStudy (babynini):

wait how did you get rid of sin and y?

OpenStudy (babynini):

wouldn't \[x=\sqrt{1-y^2}\]

OpenStudy (jojokiw3):

Yeah. I think it is.